OpenStudy (anonymous):

\frac{ 3 }{ 4 }t + \frac{ 5 }{ 6 } = \frac{ 2 }{ 3 }t

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }t + \frac{ 5 }{ 6 } = \frac{ 2 }{ 3 }t\]

OpenStudy (anonymous):

Solve for the equation and state the solution Any solution, no solution, one solution, zero solution

OpenStudy (anonymous):

@PRAETORIAN.10

OpenStudy (anonymous):

\[\frac 34 t - \frac 23 t = -\frac 56\] \[\implies t(\frac 34 - \frac 23) = -\frac 56\] \[\huge\implies t = -\frac{\frac 56}{\frac 34 - \frac 23}\] solve that and you have t

OpenStudy (anonymous):

I don't understand the thought process behind the modeling you did for this equation

OpenStudy (mathmath333):

easy way is to take LCM of denominators 4,6 and 3 and multiply the whole equation with the LCM

OpenStudy (anonymous):

So convert the denominators into 12?

OpenStudy (mathmath333):

yes absolutely

OpenStudy (anonymous):

Yeah I get 9/12t + 10/12 = 8/12t

OpenStudy (mathmath333):

now multiply and divide (10/12) with t

OpenStudy (anonymous):

I'm not sure what you mean...do 10/12 * t ?

OpenStudy (mathmath333):

all denominators except the 10/12 doesn't have t

OpenStudy (mathmath333):

\[10/12=10t/12t\]

OpenStudy (mathmath333):

like the above

OpenStudy (mathmath333):

then all the denominators will be the same and will be cancelled

OpenStudy (anonymous):

Seems kinda weird to just pull out a t and slap it on a number out of the blue, haha, maybe that's just c;

OpenStudy (mathmath333):

it saves time ,lol

OpenStudy (anonymous):

19/12t = 8/12t

OpenStudy (anonymous):

?

OpenStudy (mathmath333):

\[3t/4 +5/6=2t/3\]

OpenStudy (anonymous):

Yeah 9/12t + 10/12 = 8/12t

OpenStudy (mathmath333):

\[9t/12 +10/12=8t/12\]

OpenStudy (anonymous):

Then we multiplied 10/12 by t, to get 10/12t

OpenStudy (anonymous):

For work shown, I just did 9/12t + (t) 10/12 = 8/12t to show that I multiplied t with 10/12

OpenStudy (anonymous):

to get 10/12t

OpenStudy (mathmath333):

here if we multiply by 12 to both sies all the denominator will be canceled

OpenStudy (mathmath333):

sry

OpenStudy (mathmath333):

look at the equation now

OpenStudy (mathmath333):

\[9t+10=8t\]

OpenStudy (anonymous):

To make it 9t + 10 =8t

OpenStudy (mathmath333):

\[9t -8t+10=0\]

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

1t + 10 = 0

OpenStudy (mathmath333):

i think u can calulate {t} easily now

OpenStudy (anonymous):

or just t

OpenStudy (anonymous):

Minus 10 from both sides

OpenStudy (anonymous):

t = -10

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

Thanks so much @mathmath333 ;3

OpenStudy (mathmath333):